Calculate relative rate of effusion of SO2 to CH4 if the mixture obtained by effusing out a mixture nSO2/nCH4 =8:1 for three effusing steps

Grade:11

1 Answers

Arun
6 years ago
 
Dear Akash
 
First you must know the Rate of effusion formulae.
That is:
R = PA/root(MT)
For SO2,
R1 = P1A/root(64T)................(1)
P1V = 8nRT..............(2)
For CH4,
R2 = P2A/root(16T)...............(3)
P2V = nRT.....................(4)
divide eqn. (1) and (3)
and eqn. (2) and (4),
we get,
 
R1/R2 = (P1/P2)*root(16/64) =P1/2P2
and
P1/P2 = 8
Therfore,
R1/R2 = 8/2 = 4/1 => 4:1
is the answer. 
 
Regards
Arun (askIITians forum expert)